High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingReg 1 Reg 2 Reg 3A Reg 3B Compute the total overhead cost applied to Job 205. (Round Predetermined overhead rate to 2 decimWhat would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead rate to 2 decimal places. RoReg 1 Reg 2 Reg 3A Reg 38 If the job contained 25 units, what would be the unit product cost? (Round Predetermined overheadReg 1 Reg 2 Reg 3A Reg 3B What would be the total manufacturing cost recorded for Job 2057 (Round Predetermined overhead rat
Answer #1

1)

Predetermined overhead rate used in molding department = Estimated manufacturing overhead cost / Estimated machine hours + Variable manufacturing overhead cost per machine hour

= 221400 / 82000 + 2.80

= 2.70 + 2.80

= $5.50 per machine hours

Predetermined overhead rate used in painting department = Estimated manufacturing overhead cost / Estimated direct labor hours + Variable manufacturing overhead cost per direct labor hour

= 544320 / 56700 + 4.80

= 9.6+4.80

= $14.40 per labour hours

2)

Overhead cost applied to Job 205 = Machine hours in molding department * Overhead rate for molding department + Direct labor hours in paining department + Overhead rate for paining department

= 320 * 5.50 + 135 * 14.40 = $1760+$1944

= $3704

3 a)

Molding painting Total
Direct material $936 $1160 $2096
Direct labour $720 $970 $1690
Manufacturing overhead $1760 $1944 $3704
Total cost $3416 $4074 $7490

b)

Unit product cost = $7490/25 = 299.6